# Capillarity in Stationary Random Granular Media: Distribution-Aware Screening and Quantitative Supercell Sizing

### v2

@@ Abstract
We develop a quantitative framework to determine the minimal periodic supercell required for representative simulations of capillarity-screened Darcy flow in stationary random, polydisperse granular media. The microstructure is characterized by two-point statistics (covariance and spectral density) that govern finite-size fluctuations. Capillarity is modeled as a screened, modified-Helmholtz problem with phase-dependent transport under periodic boundary conditions; periodic homogenization yields an apparent conductivity, an apparent screening parameter, and a macroscopic capillary decay length. Because screening imparts a spatial low-pass response, we introduce a distribution-aware treatment of polydispersity consisting of a capillarity-weighted volume fraction and a screened analogue of the integral range that preserves variance units and recovers classical descriptors in the appropriate limits. These descriptors lead to two sizing rules: (i) a length criterion on the shortest cell edge controlled by a microstructural correlation length, the macroscopic decay length, and a high quantile of grain size; and (ii) a volume criterion that links the target coefficient of variation to the screened integral range and the phase contrast. The framework couples statistical microstructure information to capillary response and yields reproducible, distribution-aware supercell selection for image-based finite-element or fast-Fourier-transform solvers.
@@ I Introduction
We address this gap using the statistical framework of random heterogeneous media [Torquato2002, Chiu2013, DaleyVJ2003]. The microstructure is modeled as a stationary marked point process of grain centers and sizes; the associated two-phase indicator field yields standard two-point descriptors—the covariance and its spectral density—that control finite-size fluctuations. A single scalar summary, the integral range (a correlation volume), emerges as the key parameter governing variance with supercell volume; its definition and variance implications are reviewed in §II.2, building on [LuTorquato1990, QuintanillaTorquato1997, Torquato2002, Torquato2018]. In parallel, a low-wavenumber coverage requirement ensures that the smallest Fourier mode of the supercell falls well inside the regime that dictates finite-size bias; we formalize this in §II.3 (see also the Wiener–Khintchine correspondence recalled in §II.1, Eqs. (11b)–(11c)).
@@
On the physics side, we cast small-slope capillarity as a linear screened (modified Helmholtz/Yukawa) cell problem posed on the periodic torus, with phase-dependent transport in grains and matrix. Periodic homogenization then furnishes an apparent conductivity and an apparent screening parameter, which together define a macroscopic decay length for capillary disturbances; the formulation and cell problems are detailed in §II.6 and draw on classical capillarity and homogenization theory [DeGennes2004, LandauLifshitz1987, BensoussanLionsPapanicolaou2011, JikovKozlovOleinik1994], solved efficiently with finite-element (FE) or fast-Fourier-transform (FFT) solver methods [MoulinecSuquet1998, Geers2010].
@@
Our key observation is that screened capillarity acts as a spatial low-pass filter: structural modes with wavenumber below the inverse capillary length contribute most strongly. This motivates a distribution-aware treatment of polydispersity. We introduce a capillarity-weighted volume fraction and a screened analogue of the integral range that preserve units and reduce to their classical counterparts in the appropriate limits; these are defined precisely in §III. The resulting sizing rules are quantitative and verifiable: (i) a length criterion on the shortest cell edge that depends on a microstructural correlation length, the macroscopic capillary decay length, and a high quantile of grain size (to avoid periodic self-interaction); and (ii) a volume criterion that delivers the supercell size needed to meet a target coefficient of variation with a given number of realizations. Directional variants handle anisotropy; a posteriori spectral checks confirm adequate low-$k$ coverage [Torquato2002, Torquato2018].
@@
In summary, the paper links statistical microstructure descriptors to screened capillary physics and distills them into practical, distribution-aware rules for periodic FE/FFT simulations. It complements reconstruction and descriptor advances [YeongTorquato1998, JiaoPNAS2009] and sharpens RVE-size analyses for random composites [Kanit2003, Ostoja2006] by adding the specific screening length scale and a spectral perspective directly tied to capillarity.
@@ II.1 Problem statement and notation
where $p_{A}$ is the size density, $\nu$ is the (constant) number density of centers [Chiu2013, DaleyVJ2003].
@@
i.e., $\chi=1$ inside grains and $\chi=0$ in the matrix [Torquato2002].
@@
Note that $C$ is even, $C(-\mathbf{r})=C(\mathbf{r})$, and $C(\mathbf{0})=\phi(1-\phi)$ and the behavior of $\widehat{C}(\mathbf{k})$ as $\mathbf{k}\to\mathbf{0}$ controls the long-wavelength fluctuations that dominate finite-size effects [Torquato2002]. Moreover, the inverse relation $C(\mathbf{r})=\frac{1}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}\widehat{C}(\mathbf{k})\,e^{i\mathbf{k}\cdot\mathbf{r}}\,d\mathbf{k}$ makes the Fourier pair explicit the Wiener-Khintchine theorem [Wiener1930, Khintchine1934, Yaglom1987]: for stationary spatial random fields, the covariance and power spectral density are a Fourier transform pair. In computations we work on a periodic torus (below), where the spectrum is sampled at discrete lattice wavenumbers.
@@
where $H^{1}_{\#}(Y):=\{v\in H^{1}_{\mathrm{loc}}(\mathbb{R}^{3}):v\text{ is }Y\text{-periodic}\}$ is the periodic Sobolev space. The zero-mean condition fixes the additive constant. This periodic/affine-periodic setting is Hill–Mandel consistent and is efficiently solvable by FFT- or FE-based methods [Hill1963, Geers2010, MoulinecSuquet1998]. In later sections we use the same periodic framework to treat the screened capillarity problem.
@@ II.2 Integral range and finite-size variance
with $C(\mathbf{0})=\phi(1-\phi),\quad A>0,$ which compresses the overall strength of spatial correlations into a single volume scale (typically $A=O(\xi^{3})$ when $C$ decays beyond a correlation length $\xi$) [Torquato2002].
@@
Using stationarity and the change of variables $(\mathbf{x},\mathbf{y})\mapsto(\mathbf{x},\mathbf{r}=\mathbf{y}-\mathbf{x})$ yields the exact window identity [LuTorquato1990, QuintanillaTorquato1997, Torquato2002]:
@@
i.e., the classical $1/V$ decay of local volume-fraction variance [LuTorquato1990, QuintanillaTorquato1997, Torquato2002]. The same result appears in Fourier space by writing $W_{Y}(\mathbf{k}):=\int_{Y}e^{-i\mathbf{k}\cdot\mathbf{x}}\,d\mathbf{x}$ and using
@@
where the kernel $|W_{Y}|^{2}/V^{2}$ concentrates near $\mathbf{k}=\mathbf{0}$ as $Y$ grows, so the leading term is governed by $\widehat{C}(\mathbf{0})=\int C$ and hence by $A$ [LuTorquato1990, Torquato2002].
@@
with a finite prefactor $C_{P}=O(1)$ that depends on the physics (contrast, loading, boundary conditions) but not on $V$ [Kanit2003, Ostoja2006]. In (31), $V=L_{x}L_{y}L_{z}$ is the supercell volume; $A$ is the integral range (Eq. (19), units of volume), which measures the correlation volume of $\chi$; and $C_{P}$ is a physics-dependent prefactor with the units of $P^{2}$ (e.g., $C_{P}=C(\mathbf{0})=\phi(1-\phi)$ for $P=\widehat{\phi}$, and Eq. (80) for $P=\beta_{\rm app}$). Thus $\operatorname{Var}[P(Y)]$ decays like (correlation volume)/(sample volume), scaled by the property sensitivity. Averaging $n$ independent realizations gives
@@
which makes the dependence on $(K_{g},K_{m},\phi)$ fully transparent. For stronger contrasts or non-dilute regimes, one may replace (41) by a self-consistent estimate (e.g., Bruggeman) or compute $dP/d\phi$ numerically; the scaling $\operatorname{Var}[P(Y)]\approx(C_{P}A)/V$ remains the guiding law [Kanit2003, Ostoja2006].
Two remarks delimit the regime of validity. If $C\notin L^{1}(\mathbb{R}^{3})$ (very long-range correlations), then $A$ diverges and the $1/V$ law breaks down; variance decays more slowly and much larger cells are required [Torquato2002]. In hyperuniform media one has $\widehat{C}(\mathbf{0})=\int C=0$, hence $A=0$ and the variance decays faster than $1/V$, with exponents set by the small-$k$ behavior of $\widehat{C}$ [Torquato2018].
@@ II.3 Low-wavenumber resolution (length) constraint
If the shortest edge is too small, the window spectrum $|W_{Y}(\mathbf{k})|^{2}/V^{2}$ never probes the asymptotic low-$k$ region of $\widehat{C}$, and the $A/V$ variance law (Section II.2) is not yet observed [Torquato2002].
@@
Here $c=\mathcal{O}(10)$ is a heuristic safety factor, not a universal constant: values $c\in[8,12]$ typically place $k_{\min}$ at least a factor 551010 below the knee of $\widehat{C}(k)$ for short-ranged covariances, which in turn makes the normalized intersection factor $\alpha_{Y}(\mathbf{r})$ close to $1$ over the main support of $C$. The appropriate $c$ depends on how sharply $\widehat{C}(k)$ turns near $k=0$ and on the acceptable bias. We therefore verify a posteriori by plotting (optionally isotropized) $\widehat{C}(k)$, marking $k_{\min}=2\pi/\min(L_{\alpha})$, and confirming $k_{\min}\ll k_{c}$ for the material at hand [Torquato2002, Torquato2018]. For hyperuniform media, where $\widehat{C}(\mathbf{0})=0$ and $\widehat{C}(\mathbf{k})\sim|\mathbf{k}|^{\alpha}$ as $\mathbf{k}\to\mathbf{0}$, $L$ must be chosen so that $k_{\min}$ lies inside the power-law regime; this can require $L$ beyond the conservative $c\,\xi$ threshold [Torquato2018].
@@ II.4 Geometric safeguard for polydispersity
as stated in standard Molecular Dynamics texts and manuals [AllenTildesley2017, GromacsManual]. If the relevant interaction range scales with the large-grain size, say $R_{c}\simeq\gamma\,a_{\max}$ (with $\gamma\in[1,2]$ depending on the physics—hard/soft contact shells, lubrication, neighbor-search radius), then (53) implies
@@
which strictly contains both the non-overlap condition (52) and the minimum-image implication (54) for typical $\gamma$. For markedly anisotropic grains, replace $a_{\max}$ by a circumscribed diameter (maximum chord) for the family; if pair interactions extend beyond hard contact, use the corresponding effective $R_{c}$ when mapping to (53).
The safeguard (55) is orthogonal to the variance and low-$k$ constraints from §II.2–§II.3. The $A/V$ law sets how much volume is needed for statistical stabilization [Kanit2003, Ostoja2006, Torquato2002], and $k_{\min}\ll k_{c}$ ensures the box is long enough to probe the correct small-$k$ regime [Torquato2002, Torquato2018]. By contrast, (55) is a local, geometry-driven hygiene condition that prevents the few largest grains from interacting too strongly with their own periodic copies. Enforcing (55) avoids artifacts (image self-contact, near-contact bias, aliasing of neighbor shells) that can persist even if $V$ is large and $k_{\min}$ is small. In the same spirit, the choice $c_{a}\in[3,5]$ in Eq. (55) should be viewed as a conservative buffer: choosing $c_{a}$ too small can leave residual self-interaction of the largest grains (as seen in Sec. IV.4), whereas choosing it slightly larger only increases the box volume without altering the form of the sizing rules.
It is cheap to evaluate from the size distribution alone via (51), requires no PDE solves or spectrum estimates, and provides a quick sanity check before expensive sampling. It also matches the minimum-image requirement (53) standard in particle simulations with periodic boundaries [AllenTildesley2017, GromacsManual]. In practice, apply (55) first (cheap), then verify statistical targets with (33) and spectral reach with (47).
@@ II.5 Screened Darcy with Phase-Dependent Transport
For small interfacial slopes under gravity, the linearized Young–Laplace balance $\gamma\,\nabla^{2}h=\Delta\rho\,g\,h$ implies a screened, modified-Helmholtz operator for the interface height, $\nabla^{2}h-\lambda^{-2}h=0$ with capillary length $\lambda^{2}=\gamma/(\Delta\rho g)$. The associated Green function in 2D decays as $K_{0}(r/\lambda)$ (where $K_{0}$ is the modified Bessel function of the second kind), highlighting exponential screening. These results are classical in capillarity, see e.g. [Kralchevsky2000, DeGennes2004]. In porous media, Darcy-type models employ phase-dependent transport (mobilities/relative permeabilities), often cast via a global-pressure formulation and capillary pressure closures [Leverett1941, ChaventJaffre1986]. Together, these strands motivate a bulk equation with a Darcy operator augmented by a capillary screening term and phase-dependent transport coefficients.
@@
where $\lambda=\sqrt{\gamma/(\Delta\rho g)}$ encodes capillarity (surface tension $\gamma$, density contrast $\Delta\rho$, gravity $g$), and $\rho_{0}$ collects injections/ejections. In the limit $\lambda\to\infty$, Eq. (56) reduces to the classical Darcy/Poisson form. With constant $K$, Eq. (56) is the modified Helmholtz (Yukawa) equation. Numerical aspects of the modified Helmholtz operator and its Green-function structure are discussed in [Winkelmann2021].

### v3

@@ Abstract
We develop a quantitative framework to determine the minimal periodic supercell required for representative simulations of capillarity-screened Darcy flow in stationary random, polydisperse granular media. The microstructure is characterized by two-point statistics (covariance and spectral density) that govern finite-size fluctuations. Capillarity is modeled as a screened, modified-Helmholtz problem with phase-dependent transport under periodic boundary conditions; periodic homogenization yields an apparent conductivity, an apparent screening parameter, and a macroscopic capillary decay length. Because screening imparts a spatial low-pass response, we introduce a distribution-aware treatment of polydispersity consisting of a capillarity-weighted volume fraction and a screened analogue of the integral range that preserves variance units and recovers classical descriptors in the appropriate limits. These descriptors lead to two sizing rules: (i) a length criterion on the shortest cell edge controlled by a microstructural correlation length, the macroscopic decay length, and a high quantile of grain size; and (ii) a volume criterion that links the target coefficient of variation to the screened integral range and the phase contrast. The framework couples statistical microstructure information to capillary response and yields reproducible, distribution-aware supercell selection for image-based finite-element or fast-Fourier-transform solvers. The resulting criteria are therefore intended for representativity of the coarse-grained screened response, rather than for isolated nonlinear pore-scale events.
@@ I Introduction
We address this gap using the statistical framework of random heterogeneous media [22, 5, 6]. The microstructure is modeled as a stationary marked point process of grain centers and sizes; the associated two-phase indicator field yields standard two-point descriptors—the covariance and its spectral density—that control finite-size fluctuations. A single scalar summary, the integral range (a correlation volume), emerges as the key parameter governing variance with supercell volume; its definition and variance implications are reviewed in §II.2, building on [18, 21, 22, 23]. In parallel, a low-wavenumber coverage requirement ensures that the smallest Fourier mode of the supercell falls well inside the regime that dictates finite-size bias; we formalize this in §II.3 (see also the Wiener–Khintchine correspondence recalled in §II.1, Eqs. (11b)–(11c)).
@@
On the physics side, we cast small-slope capillarity as a linear screened (modified Helmholtz/Yukawa) cell problem posed on the periodic torus, with phase-dependent transport in grains and matrix. Periodic homogenization then furnishes an apparent conductivity and an apparent screening parameter, which together define a macroscopic decay length for capillary disturbances; the formulation and cell problems are detailed in §II.6 and draw on classical capillarity and homogenization theory [7, 16, 3, 12], solved efficiently with finite-element (FE) or fast-Fourier-transform (FFT) solver methods [19, 8].
@@
Our key observation is that screened capillarity acts as a spatial low-pass filter: structural modes with wavenumber below the inverse capillary length contribute most strongly. This motivates a distribution-aware treatment of polydispersity. We introduce a capillarity-weighted volume fraction and a screened analogue of the integral range that preserve units and reduce to their classical counterparts in the appropriate limits; these are defined precisely in §III. The resulting sizing rules are quantitative and verifiable: (i) a length criterion on the shortest cell edge that depends on a microstructural correlation length, the macroscopic capillary decay length, and a high quantile of grain size (to avoid periodic self-interaction); and (ii) a volume criterion that delivers the supercell size needed to meet a target coefficient of variation with a given number of realizations. Directional variants handle anisotropy; a posteriori spectral checks confirm adequate low-$k$ coverage [22, 23].
@@
In summary, the paper links statistical microstructure descriptors to screened capillary physics and distills them into practical, distribution-aware rules for periodic FE/FFT simulations. It complements reconstruction and descriptor advances [29, 11] and sharpens RVE-size analyses for random composites [13, 20] by adding the specific screening length scale and a spectral perspective directly tied to capillarity.
@@ II.1 Problem statement and notation
where $p_{A}$ is the size density, $\nu$ is the (constant) number density of centers [5, 6]. The factorization in Eq. (4) is a convenient baseline, not a structural requirement of the sizing theory. In dense or jammed packings, steric exclusion and contact networks generally induce correlations between positions and marks, so the independent-mark assumption can fail. The supercell criteria developed below remain applicable provided the relevant two-point descriptors $C(\mathbf{r})$, $\widehat{C}(\mathbf{k})$, $A$, or their screened counterparts are estimated directly from the actual microstructure (experimental image, DEM packing, or correlated generator). In that case, Eqs. (4) are replaced by the measured statistics of the correlated medium rather than assumed a priori.
@@
i.e., $\chi=1$ inside grains and $\chi=0$ in the matrix [22].
@@
Note that $C$ is even, $C(-\mathbf{r})=C(\mathbf{r})$, and $C(\mathbf{0})=\phi(1-\phi)$ and the behavior of $\widehat{C}(\mathbf{k})$ as $\mathbf{k}\to\mathbf{0}$ controls the long-wavelength fluctuations that dominate finite-size effects [22]. Moreover, the inverse relation $C(\mathbf{r})=\frac{1}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}\widehat{C}(\mathbf{k})\,e^{i\mathbf{k}\cdot\mathbf{r}}\,d\mathbf{k}$ makes the Fourier pair explicit the Wiener-Khintchine theorem [25, 14, 28]: for stationary spatial random fields, the covariance and power spectral density are a Fourier transform pair. In computations we work on a periodic torus (below), where the spectrum is sampled at discrete lattice wavenumbers.
@@
where $H^{1}_{\#}(Y):=\{v\in H^{1}_{\mathrm{loc}}(\mathbb{R}^{3}):v\text{ is }Y\text{-periodic}\}$ is the periodic Sobolev space. The zero-mean condition fixes the additive constant. This periodic/affine-periodic setting is Hill–Mandel consistent and is efficiently solvable by FFT- or FE-based methods [10, 8, 19]. In later sections we use the same periodic framework to treat the screened capillarity problem.
@@ II.2 Integral range and finite-size variance
with $C(\mathbf{0})=\phi(1-\phi),\quad A>0,$ which compresses the overall strength of spatial correlations into a single volume scale (typically $A=O(\xi^{3})$ when $C$ decays beyond a correlation length $\xi$) [22].
@@
Using stationarity and the change of variables $(\mathbf{x},\mathbf{y})\mapsto(\mathbf{x},\mathbf{r}=\mathbf{y}-\mathbf{x})$ yields the exact window identity [18, 21, 22]:
@@
i.e., the classical $1/V$ decay of local volume-fraction variance [18, 21, 22]. The same result appears in Fourier space by writing $W_{Y}(\mathbf{k}):=\int_{Y}e^{-i\mathbf{k}\cdot\mathbf{x}}\,d\mathbf{x}$ and using
@@
where the kernel $|W_{Y}|^{2}/V^{2}$ concentrates near $\mathbf{k}=\mathbf{0}$ as $Y$ grows, so the leading term is governed by $\widehat{C}(\mathbf{0})=\int C$ and hence by $A$ [18, 22].
@@
with a finite prefactor $C_{P}=O(1)$ that depends on the physics (contrast, loading, boundary conditions) but not on $V$ [13, 20]. In (31), $V=L_{x}L_{y}L_{z}$ is the supercell volume; $A$ is the integral range (Eq. (19), units of volume), which measures the correlation volume of $\chi$; and $C_{P}$ is a physics-dependent prefactor with the units of $P^{2}$ (e.g., $C_{P}=C(\mathbf{0})=\phi(1-\phi)$ for $P=\widehat{\phi}$, and Eq. (80) for $P=\beta_{\rm app}$). Thus $\operatorname{Var}[P(Y)]$ decays like (correlation volume)/(sample volume), scaled by the property sensitivity. Averaging $n$ independent realizations gives
@@
which makes the dependence on $(K_{g},K_{m},\phi)$ fully transparent. For stronger contrasts or non-dilute regimes, one may replace (41) by a self-consistent estimate (e.g., Bruggeman) or compute $dP/d\phi$ numerically; the scaling $\operatorname{Var}[P(Y)]\approx(C_{P}A)/V$ remains the guiding law [13, 20].
Two remarks delimit the regime of validity. If $C\notin L^{1}(\mathbb{R}^{3})$ (very long-range correlations), then $A$ diverges and the $1/V$ law breaks down; variance decays more slowly and much larger cells are required [22]. In hyperuniform media one has $\widehat{C}(\mathbf{0})=\int C=0$, hence $A=0$ and the variance decays faster than $1/V$, with exponents set by the small-$k$ behavior of $\widehat{C}$ [23].
@@ II.3 Low-wavenumber resolution (length) constraint
If the shortest edge is too small, the window spectrum $|W_{Y}(\mathbf{k})|^{2}/V^{2}$ never probes the asymptotic low-$k$ region of $\widehat{C}$, and the $A/V$ variance law (Section II.2) is not yet observed [22].
@@
Here $c=\mathcal{O}(10)$ is a heuristic safety factor, not a universal constant: values $c\in[8,12]$ typically place $k_{\min}$ at least a factor 551010 below the knee of $\widehat{C}(k)$ for short-ranged covariances, which in turn makes the normalized intersection factor $\alpha_{Y}(\mathbf{r})$ close to $1$ over the main support of $C$. The appropriate $c$ depends on how sharply $\widehat{C}(k)$ turns near $k=0$ and on the acceptable bias. We therefore verify a posteriori by plotting (optionally isotropized) $\widehat{C}(k)$, marking $k_{\min}=2\pi/\min(L_{\alpha})$, and confirming $k_{\min}\ll k_{c}$ for the material at hand [22, 23]. For hyperuniform media, where $\widehat{C}(\mathbf{0})=0$ and $\widehat{C}(\mathbf{k})\sim|\mathbf{k}|^{\alpha}$ as $\mathbf{k}\to\mathbf{0}$, $L$ must be chosen so that $k_{\min}$ lies inside the power-law regime; this can require $L$ beyond the conservative $c\,\xi$ threshold [23].
@@ II.4 Geometric safeguard for polydispersity
as stated in standard Molecular Dynamics texts and manuals [1, 9]. If the relevant interaction range scales with the large-grain size, say $R_{c}\simeq\gamma\,a_{\max}$ (with $\gamma\in[1,2]$ depending on the physics—hard/soft contact shells, lubrication, neighbor-search radius), then (53) implies
@@
which strictly contains both the non-overlap condition (52) and the minimum-image implication (54) for typical $\gamma$. For markedly anisotropic grains (fibers, plates, elongated clasts), $a_{\max}$ should be interpreted as the upper quantile of a geometric extent relevant to periodic self-interaction, not as an equivalent-sphere diameter. A robust choice is the maximum chord length, or equivalently the diameter of the circumscribed sphere, for the grain family:
$a_{\max}^{\rm geom}:=Q_{1-\delta}(d_{\rm circ}).$
For highly elongated particles, near-parallel image interactions can persist over distances larger than the minor axis, so the safety factor $c_{a}$ may need to be chosen toward the upper end of, or modestly above, the range $[3,5]$. If the medium is directionally anisotropic, a directional safeguard based on projected extents along each axis may be used instead.
The safeguard (55) is orthogonal to the variance and low-$k$ constraints from §II.2–§II.3. The $A/V$ law sets how much volume is needed for statistical stabilization [13, 20, 22], and $k_{\min}\ll k_{c}$ ensures the box is long enough to probe the correct small-$k$ regime [22, 23]. By contrast, (55) is a local, geometry-driven hygiene condition that prevents the few largest grains from interacting too strongly with their own periodic copies. Enforcing (55) avoids artifacts (image self-contact, near-contact bias, aliasing of neighbor shells) that can persist even if $V$ is large and $k_{\min}$ is small. In the same spirit, the choice $c_{a}\in[3,5]$ in Eq. (55) should be viewed as a conservative buffer: choosing $c_{a}$ too small can leave residual self-interaction of the largest grains (as seen in Sec. IV.4), whereas choosing it slightly larger only increases the box volume without altering the form of the sizing rules.
It is cheap to evaluate from the size distribution alone via (51), requires no PDE solves or spectrum estimates, and provides a quick sanity check before expensive sampling. It also matches the minimum-image requirement (53) standard in particle simulations with periodic boundaries [1, 9]. In practice, apply (55) first (cheap), then verify statistical targets with (33) and spectral reach with (47).
@@ II.5 Screened Darcy with Phase-Dependent Transport
For small interfacial slopes under gravity, the linearized Young–Laplace balance $\gamma\,\nabla^{2}h=\Delta\rho\,g\,h$ implies a screened, modified-Helmholtz operator for the interface height, $\nabla^{2}h-\lambda^{-2}h=0$ with capillary length $\lambda^{2}=\gamma/(\Delta\rho g)$. The associated Green function in 2D decays as $K_{0}(r/\lambda)$ (where $K_{0}$ is the modified Bessel function of the second kind), highlighting exponential screening. These results are classical in capillarity, see e.g. [15, 7]. In porous media, Darcy-type models employ phase-dependent transport (mobilities/relative permeabilities), often cast via a global-pressure formulation and capillary pressure closures [17, 4]. Together, these strands motivate a bulk equation with a Darcy operator augmented by a capillary screening term and phase-dependent transport coefficients.
@@
where $\lambda=\sqrt{\gamma/(\Delta\rho g)}$ encodes capillarity (surface tension $\gamma$, density contrast $\Delta\rho$, gravity $g$), and $\rho_{0}$ collects injections/ejections. In the limit $\lambda\to\infty$, Eq. (56) reduces to the classical Darcy/Poisson form. With constant $K$, Eq. (56) is the modified Helmholtz (Yukawa) equation. Numerical aspects of the modified Helmholtz operator and its Green-function structure are discussed in [27].
We stress that this screened formulation targets the macroscopic capillary field in the small-slope regime. It is not intended to resolve strongly nonlinear pore-scale events such as snap-off, burst dynamics, or Haines jumps, whose onset depends on local throat geometry, network connectivity, and dynamic contact-line effects beyond the present linearized homogenized description. In particular, flowrate-controlled invasion percolation [26] and pressure-controlled invasion with trapping [2] exhibit scale-free (power-law) cluster statistics and long-range connectivity, which can shift REV sizing towards invasion correlation lengths and finite-size scaling rather than a finite correlation volume alone. In practice, when invasion criticality is the target observable (e.g., cluster-size or burst-size exponents, trapped-phase topology), REV sizing should instead be based on the corresponding correlation length (or finite-size scaling) of the invasion process, which can grow to the system size near threshold.